OpenStudy (anonymous):

How do I factor the following: 27c^3 -64

OpenStudy (mertsj):

Use the difference of cubes formula which is:

OpenStudy (mertsj):

\[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

Mind if I help our Mertsj?

OpenStudy (anonymous):

*out

OpenStudy (mertsj):

As long as you don't just write down the answer.

OpenStudy (anonymous):

Well, it'll be kind of difficult to not just write down the answer since it's so simple, so I guess I'll just let you handle it :).

OpenStudy (homeworksucks):

so cube root the first number, that gives you your a value, cube rooting the second gives you your b-value, and just plug those in

OpenStudy (homeworksucks):

remember subtracting a negative is adding a positive

OpenStudy (anonymous):

so would 'a' end up equaling '3c?'

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

I did the formulae, and it became a binomial multiplied by a trinomial, but when I simplified that I did not get a answers that I think is correct

OpenStudy (mertsj):

\[27c^3-64=(3c-4)(9c^2+12c+16)\]

OpenStudy (anonymous):

wouldn't that be a negative 12c?

OpenStudy (mertsj):

Did you see my previous post: a^3−b^3=(a−b)(a^2+ab+b^2)

OpenStudy (anonymous):

o yes I get it now

OpenStudy (mertsj):

Good for you.

OpenStudy (anonymous):

so is that fully simplified?

OpenStudy (anonymous):

o wait it is not

OpenStudy (mertsj):

yes it is fully simplified.

OpenStudy (anonymous):

so the fully simplified answer would be (3c-4)(3c+12)(3c-8)?

OpenStudy (mertsj):

That would be correct IF (3c+12)(3c-8)=9c^2+12c+16 Does it?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

What are you smokin' ?

OpenStudy (anonymous):

oooo... it does not.

OpenStudy (mertsj):

The factors of 27c^3-64 are (3c-4)(9c^2+12c+16) Now could we please put this problem to bed?